OpenStudy (anonymous):

Solve for x. x^x(lnx+1)=0

OpenStudy (australopithecus):

there is no Real solution to this problem

OpenStudy (australopithecus):

unless that is \[x^x(lnx + 1) = 0\]

OpenStudy (australopithecus):

x^x(lnx+1)=0 (ln(x) + 1) = 0 ln(x) = -1 x = e^(-1)

OpenStudy (australopithecus):

I divided away x^x

OpenStudy (australopithecus):

as x^x =/= 0

OpenStudy (anonymous):

okay gotcha. I was wondering where the x^x went to. Thank you!